OpenStudy (anonymous):

how do I factor y=x^2+9x+6?

OpenStudy (king):

use quadratic equation formula

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but then there is no real root

OpenStudy (king):

wait lemme try

OpenStudy (king):

\[(-9+\sqrt{57})\div2\]

OpenStudy (anonymous):

yeah that's what I got

OpenStudy (king):

thats real ...

OpenStudy (agreene):

This can actually be factored without complex numbers--you just have to think about it like you're completing a square.

OpenStudy (agreene):

\[\frac12(-2x+\sqrt{57}-9)(2x+\sqrt{57}+9)\]

OpenStudy (anonymous):

Yeah sorry you're right.